Yosad930:

;-; exam is next tuesday, and i dont get a single thing (well maybe one or two) can someone helppp

Yosad930:

SsTDOGRednek:

Yeah, I can try to help. Let me teach myself first lol

Yosad930:

@sstdogrednek wrote:
Yeah, I can try to help. Let me teach myself first lol
lol alrighty gudluck dont rush urself either, u got a few days

SsTDOGRednek:

Oh, alr then!

SsTDOGRednek:

So I google imaged it... it said Simplified expressions are 1, 2 csc x, -1 and cos ^2x. But that doesn't make sense... for the first question are you supposed to simplify all of them?

SsTDOGRednek:

Oh wait... Simplify each trigonometric expression using trigonometric identities.

SsTDOGRednek:

Step 1: 1+cot^2x=1/sin^2x sin^2x(1+cot^2x) = sin^2x times 1/siin^2x sin^2x(1+cot^2x)=1 Step 2: Find a common denominator and add the fractions. 1-cosx/sinx + sinx/1-cosx = (1-cos x)^2 +sin^2x/sin x(1-cosx) Expand and simplify the numerator: I wish I didn't start doing this... (1-2cosx+cos^2x)+sin^2 x/ sinx(1-cosx) 1-2csx+(cos^2x+sin^2x)/sinx(1-cosx) Use the identity sin^2x+cos^2x=1: 1-2cosx+1/sinx(1-cosx)=2-2cosx/sinx(1-cosx) Factor and simplify: 2(1-cosx)/sinx(1-cosx)= 2/sin x Use the identity csc x= 1/sinx: 2/sinx=2cscx Step 3: Use the identity 1+tan^2x=sec^2x: tan^2x-sec^2c=tan^2x-(1+tan^2x) step 4: Simplify: 2+tan^2x/sec^2x - 1 Solution: simplified solutions are: 1,2 cscx, -1 and cos^2x I hate myself now... lol

SsTDOGRednek:

I really hope that helps because I spent like too much time on that...

SsTDOGRednek:

Shoot... that was only the first question... NOOOOOOOOOOOOOOOOOOOOOOOOOOOOOO!!!

SsTDOGRednek:

You might have to use PEMDAS in some of the problems, but idk though...

KnoxxyBoy:

ur scaring me with numbers...

SsTDOGRednek:

I'M GOING CRAZY!!!

Yosad930:

srry im backk lol yeah ik im going crazy with this too, and i got taught by an actual college professor who only talks and talks and talks

SsTDOGRednek:

Sounds like my geometry teacher, lol

Yosad930:

@sstdogrednek wrote:
Sounds like my geometry teacher, lol
yea imma look at it later.. dealing with problems after problems atm..